IfX~ U(37,84), then P(X> 46|X<71) = 0.5319 0.8085 0.6923 0.7353
Answer #1

Solution-

If X ~ U (37, 84)

(Uniformily distributed between 37 and 84).

So, Probability density function is

f(X) = 1/(84-37) = 47

Now, the required probability is

P(X > 46| X < 71)

= P( 46< X < 71)/P(37 <X<71)

= [(71-46)/f(x)] / [(71-37)/f(x)]

= (71 -46)/(71-37)

= 25/34

=0.735291176

≈0.7353

Hence, P(X>46 | X <71) = 0.7353